# Bridging Gaps, Building Futures: Advancing Software Developer Diversity and Inclusion Through Future-Oriented Research

### v1

@@ Abstract.
Software systems are responsible for nearly all aspects of modern life and society. However, the demographics of software development teams that are tasked with designing and maintaining these software systems rarely match the demographics of users. As the landscape of software engineering (SE) evolves due to technological innovations, such as the rise of automated programming assistants powered by artificial intelligence (AI) and machine learning, more effort is needed to promote software developer diversity and inclusion (SDDI) to ensure inclusive work environments for development teams and usable software for diverse populations. To this end, we present insights from SE researchers and practitioners on challenges and solutions regarding diversity and inclusion in SE. Based on these findings, we share potential utopian and dystopian visions of the future and provide future research directions and implications for academia and industry to promote SDDI in the age of AI-driven SE.
@@
$c$s: Software and its engineering Programming teamsccs: Human-centered computing Accessibility theory, concepts and paradigmsccs: Applied computing Educationccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
Software systems are ubiquitous in society and impact nearly all aspects of modern life. Software engineering (SE), that is, the processes, methods, and tools to support the development and maintenance of software (Pressman, 2005), is crucial for producing high-quality applications that impact human behavior, well-being, and decision-making. Recent innovations–such as the advent of large language models (LLMs) and machine learning-based systems–have transformed the software development landscape and introduced novel approaches to automate and support SE tasks (de Souza et al., 2024; White et al., 2023). For example, GitHub states that, as of February 2023, their coding assistant Copilot is “behind an average of 46% of developers’ code across all programming languages.” [^1] https://github.blog/2023-02-14-github-copilot-now-has-a-better-ai-model-and-new-capabilities/
@@
Despite recent technological innovations, SE is largely dependent on the efforts of software developers as “significant human involvement and expertise” (White et al., 2023, p. 13) are necessary to leverage LLM to automate tasks related to the design, implementation, testing, and maintenance of applications. However, the diversity and values of software development teams that design software systems often do not reflect the diversity and values of intended users, or, more broadly, our society. For example, in the 2022 Stack Overflow Developer Survey, approximately 92% of the respondents identified as male and 77% identified as White or European (Overflow, [n. d.]). In contrast, the global population is approximately 50% male (Dyvik, 2023) and 16% White (Torre, 2018).
@@
This “diversity crisis” (Albusays et al., 2021) in software development can have major ramifications for people from underrepresented backgrounds. For example, research shows that most software lacks gender inclusivity, favoring the problem-solving processes of men (Burnett et al., 2016a). Moreover, this crisis contributes to non-inclusive environments where minority developers are disadvantaged. For example, peer code reviews, where developers review code from contributors before merging into source code, are a common practice to improve software quality (McIntosh et al., 2016). However, recent studies show that code contributions from programmers of non-White and non-male backgrounds receive more pushback (Murphy-Hill et al., 2022) and higher rejection rates (Nadri et al., 2021; Terrell et al., 2017).
@@
Research reveals that diversity and inclusion enhance software development (Dowler, 2023; (2019), NSB). For example, studies show that open source projects with diverse contributors are more productive (Vasilescu et al., 2015), neuroinclusive teams are more productive than purely neurotypical teams (Austin and Pisano, 2017), heterogeneous collaboration based on race and ethnicity leads to a higher number of contributions to open source projects (Shameer et al., 2023), and working on gender-diverse teams improves attitudes towards women and improves decision-making and innovation (Kohl and Prikladnicki, 2022). However, a significant amount of work must be done to create inclusive work environments that lead to a more diverse community, which is building the software that is the foundation of our digital society.
To this end, this paper shares insights from SE researchers and practitioners on the challenges and opportunities with respect to diversity and inclusion. We begin by sketching four contrasting scenarios for SE in 2030, two outlining a utopian future, and two outlining a dystopian future. We then proceed to introduce four themes that guide the discussion in the remainder of this paper: methodologies and metrics, intersectionality, knowledge transfer, and socioeconomic understanding in SE. Based on these themes, we offer guidance for researchers to bridge the gaps in software developer diversity and inclusion (SDDI). Finally, we discuss how the rapid developments in artificial intelligence (AI) can challenge or support progress in SDDI, and how a carefully balanced use of AI is essential for building an inclusive future that avoids dystopian scenarios and gets as close as possible to a utopian future.
@@ 2. Research Process
The results presented in this paper originate in discussions during an academic meeting in June 2023.222https://shonan.nii.ac.jp/seminars/194 This meeting brought together 23 software engineering researchers and practitioners from diverse backgrounds and career phases interested in fostering SDDI-related research. 18 workshop participants were actively conducting SE research at academic institutions from around the world, while five were researchers and/or developers at companies in industry. The workshop was organized in a hybrid format, with seven participants joining remotely.
@@
To foster inclusive participation and generate a wide range of ideas from participants, we employed the liberating structure known as 1-2-4-ALL (Lipmanowicz and McCandless, 2013). [^3] https://www.liberatingstructures.com/1-1-2-4-all/ Liberating structures are qualitative discussion techniques to support lively discussions and foster engagement in a group setting (lib, [n. d.]). We utilized 1-2-4-ALL as a data collection method to allow self-reflection and collaborative discussion, building toward consensus or shared understanding among participants. Participants were divided into three breakout groups and tasked with producing one goal, two outcomes, and four themes related to challenges and opportunities to improve SDDI. The results of the breakout group discussions were brought to the broader group of participants to obtain feedback and reach agreement on a defined list of SDDI themes. The participants were then divided into groups according to the defined themes to further expand on the research challenges and solutions. After intensive discussions around the identified goals, outcomes, and themes in multiple group sessions, we developed a report [^4] https://shonan.nii.ac.jp/docs/No%20.194.pdf summarizing the discussion for each theme.
@@
In this paper, we expand on four of these themes with the aim of building a more coherent vision of future research directions for SDDI: methodologies and metrics, intersectionality, large-scale socioeconomic data, and knowledge transfer. These themes were chosen based on the timeliness, potential impact, and interest participants attributed to the themes. From each of the themes, we identified potential benefits and harms. In turn, the themes and related research literature guided us to form research goals for 2030.
@@ 3. Four Scenarios for 2030
In the following, we present two utopian scenarios, that is, scenarios “having the characteristics and organization of a perfect society” [^5] https://dictionary.cambridge.org/dictionary/english/utopian and two contrasting dystopian scenarios. These scenarios capture the future of software development work and the future of education, two areas that we, as SE researchers, can actively study and shape.
S1: Workplace Utopia: Jamie works as a recently-graduated junior software engineer at an up-and-coming tech company. The company is at the forefront of inclusion, offering, e.g., flexible work arrangements that consider factors such as child care responsibilities, support for various career paths and the age of employees and applicants, and preferences concerning work and communication modes. These initiatives are paying off, as the company has successfully attracted a diverse workforce, leading to higher productivity, a welcoming and safe work environment, and more innovative and successful products. Despite coming from a minority group and being relatively junior, Jamie feels that the company provides a safe space for developing their career and competencies and invites them to participate in decision-making. For instance, the company offers automated tools to support various workflows, e.g., in code review and in programming, that allow Jamie to receive early feedback on their work, spot mistakes, and improve their skills without fear of discrimination. Similarly, the tools and practices at the company allow diverse teams to communicate effectively and positively. The company fosters a culture that encourages the active use and development of these tools for everyone’s benefit, both within and beyond the company. In addition, the company carefully monitors existing diversity metrics and plans interventions based on them. Regular surveys help company leadership to considers employees’ values in their strategic decisions.
@@
S2: Education Utopia: Kris is a university-level educator in software engineering. Due to inclusive and equitable conditions at the university, students from varied backgrounds are enrolled in Kris’ courses, contributing to a truly diverse and inclusive study experience. Although classes are large so that the university can cope with the continuing strong demand for software engineering professionals, technology advances and government incentives make it possible for Kris to provide targeted feedback to each and every student, regardless of diversity aspects such as gender, nationality, or age. Tools and indicators are also enabling automated feedback and coaching so that Kris and their staff can fully focus on providing the best possible learning experience. Finally, the assessment has changed from standardized and unified exams and assignments to a personalized form, tailored to suit each student’s individual needs. In this environment, students feel safe to make mistakes and express themselves. In turn, this allows fruitful collaborations between students and staff.
To contrast the experiences of Jamie and Kris, we outline a dystopian future in which Ash struggles in their workplace, and Moss struggles in their role as a student.
S3: Workplace Dystopia: Ash has graduated from a prestigious university in their home country with a degree in software engineering. They are employed at a local IT company. The company provides software development services to large corporations in rich and highly developed countries. Due to the economic realities in their home country, Ash cannot make use of existing top-of-the-shelf tools, such as coding coaches or LLM-based code generation tools. This results in more manual work and less time to develop skills. Similarly, language barriers prevent them from learning and improving their skills in the same pace as similarly qualified graduates in richer and more developed countries, as existing tools cater only to English speakers. At Ash’s employer, traditional gender role models and hierarchical structures persist and affect career progression chances. As a result, Ash feels that their chance to a successful career is relatively low due to factors outside of their control. An overall lack of awareness of diversity matters also affects the services provided by Ash’s employer, as stereotypes and biases are ingrained in the labeled data.
@@
S4: Education Dystopia: Moss is a student in software engineering at a local university. Many tasks and activities that used to involve human interaction have been replaced by automated tools. This has made education cheaper for the university. However, instead of lowering tuition fees, student numbers have increased dramatically, but staff have been reduced. As a result, Moss does not feel like they belong to the university, having almost no interaction with students and staff who come from a similar background as them. As someone who moved to the city for their studies, social integration has therefore been lacking. Instead, Moss feels that education could just as well be remote. Lecture topics and assignments lack personalization and typically cover generic and stereotypical examples.
4. Results: Avoiding Dystopia

### v2

# Making Software Development More Diverse and Inclusive: Key Themes, Challenges, and Future Directions

@@ Abstract.
Introduction: Digital products increasingly reshape industries, influencing human behavior and decision-making. However, the software development teams developing these systems often lack diversity, which may lead to designs that overlook the needs, equal treatment or safety of diverse user groups. These risks highlight the need for fostering diversity and inclusion in software development to create safer, more equitable technology.
Method: This research is based on insights from an academic meeting in June 2023 involving 23 software engineering researchers and practitioners. We used the collaborative discussion method 1-2-4-ALL as a systematic research approach and identified six themes around the theme “challenges and opportunities to improve Software Developer Diversity and Inclusion (SDDI)”. We identified benefits, harms, and future research directions for the four main themes. Then, we discuss the remaining two themes, Artificial Intelligence & SDDI and AI & Computer Science education, which have a cross-cutting effect on the other themes.
Results: This research explores the key challenges and research opportunities for promoting SDDI, providing a roadmap to guide both researchers and practitioners. We underline that research around SDDI requires a constant focus on maximizing benefits while minimizing harms, especially to vulnerable groups. As a research community, we must strike this balance in a responsible way.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06journal: TOSEMjournalvolume: 0journalnumber: 0article: 0ccs: Software and its engineering Programming teamsccs: Human-centered computing Accessibility theory, concepts and paradigmsccs: Applied computing Educationccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
Software systems are ubiquitous in society, impacting nearly all aspects of modern life as a growing number of companies and industries across domains rely on and deliver products as digital services (Andreessen, 2011). For instance, recent studies show approximately 5.45 billion people, 67.1% of the global population, rely on the internet—representing a “supermajority” of the planet and still rapidly increasing around the world (Kepios, 2024). The increasing dependence on software has led to the digitization of agriculture (Bőgel, 2015), food sourcing (Paavola and Sele, 2020), environmental monitoring (Hino et al., 2018), legal proceedings (Zhurkina et al., 2021), and many other traditionally non-digitized domains. Software engineering (SE), that is, the processes, methods, and tools to support the development and maintenance of software (Pressman, 2005), is crucial for producing high-quality applications that increasingly impact human behavior, well-being, and decision-making. Recent innovations–such as the advent of generative artificial intelligence (AI) and large language models (LLMs) and machine learning-based systems–have transformed the software development landscape and introduced novel approaches to automate and support SE tasks (de Souza et al., 2024; White et al., 2023). For example, GitHub states that, as of February 2023, their coding assistant Copilot is “behind an average of 46% of developers’ code across all programming languages” [^1] https://github.blog/2023-02-14-github-copilot-now-has-a-better-ai-model-and-new-capabilities/.
@@
Generative AI is experiencing a global surge in software development activities, with the number of relevant open source projects on GitHub more than doubling in 2022—increasing by 248% with individual contributors to these projects up 148% around the world (Daigle and GitHub Staff, 2023). However, despite recent technological innovations, SE is largely dependent on the efforts of software developers. For instance, prior work concludes “significant human involvement and expertise” (White et al., 2023, p. 13) are necessary for leveraging LLMs to automate tasks related to the design, implementation, testing, and maintenance of applications, and there is a call for generative AI in software engineering to be human-centered (Russo et al., 2024). However, the diversity and values of software development teams that design software systems often do not reflect the diversity and values of intended users, or, more broadly, our society. For example, the European Institute for Gender Equality estimates that only 17% of ICT specialists in Europe are women [^2] https://eige.europa.eu/publications-resources/toolkits-guides/work-life-balance/women-in-ict. Similarly, in the 2022 Stack Overflow Developer Survey, approximately 92% of the respondents identified as male and 77% identified as White or European (Overflow, [n. d.]) [^3] More recent versions of this survey, deployed in 2023 and 2024, exclude these demographic data.. In contrast, the global population is approximately 50% male (Dyvik, 2023) and 16% White (Torre, 2018).
@@
This “diversity crisis” (Albusays et al., 2021) in SE can have major ramifications for people from underrepresented backgrounds. For example, research shows that most software lacks gender inclusivity, favoring the problem-solving processes of men (Burnett et al., 2016a). Moreover, this crisis contributes to non-inclusive environments where minority developers are disadvantaged. For example, peer code reviews, where developers review code from contributors before merging into source code, are a common practice to improve software quality (McIntosh et al., 2016). However, recent studies show that code contributions from programmers of non-White and non-male backgrounds receive more pushback (Murphy-Hill et al., 2022) and higher rejection rates (Nadri et al., 2021; Terrell et al., 2017).
@@
Research reveals that diversity and inclusion enhance SE (Dowler, 2023; (2019), NSB). For example, studies show that open source projects with diverse contributors are more productive (Vasilescu et al., 2015), neuroinclusive teams are more productive than purely neurotypical teams (Austin and Pisano, 2017), heterogeneous collaboration based on race and ethnicity leads to a higher number of contributions to open source projects (Shameer et al., 2023), and working on gender-diverse teams improves attitudes towards women and improves decision-making and innovation (Kohl and Prikladnicki, 2022). However, a significant amount of work must be done to create inclusive work environments that lead to a more diverse community, which is building the software that is the foundation of our digital society.
However, recently there have been reports of many companies scaling back in their DEI (Diversity, Equity and Inclusion) activities or teams (Burnett and Aguinis, 2024; Alfonseca and Zahn, 2023). Some organizations promoting DEI in the tech industry, such as Girls in Tech and Women Who Code, have also even been forced to shut down due to reduced financial support (Nix et al., 2024). While backlash and critical discussions around DEI initiatives have gained significant attention, especially in the media, there are also valid concerns about improving their effectiveness, accountability, and long-term impact (Follmer et al., 2024; Robb and Rana, 2024; Cox and Nguyen, 2024). At the same time, evidence supporting the business case and the importance of DEI initiatives is growing stronger (McKinsey & Company, 2023a).
Hence, there is a clear mismatch between the existing evidence that finds various benefits of diversity and inclusion in SE, and the work highlighting the rampant diversity crisis in the same field. This mismatch suggests that research findings that underline the benefits of diversity and inclusion might not be strong enough or might not be successfully transferred to industrial practice and that a more detailed understanding of the various factors contributing to the diversity crisis in SE is necessary.
To this end, this paper shares insights from SE researchers and practitioners on the challenges and opportunities regarding diversity and inclusion and creates a roadmap for avoiding undesirable outcomes (“dystopia”) and working towards a desirable state (“utopia”). Our overarching research question is:
RQ: What are the key challenges and research opportunities for promoting Software Developer Diversity and Inclusion (SDDI)?
We begin by sketching four contrasting scenarios for SE in 2030: two outlining a desirable state (“utopian future”) and two outlining an undesirable state (“dystopian future”). For these scenarios, we used six themes systematically developed from an academic meeting held in June 2023. We then proceeded to introduce the four main themes that guided the results section of this paper and stood out as clear, separate themes: methodologies and metrics, intersectionality, knowledge transfer, and socioeconomic understanding in SE. Based on these themes, we offer guidance for researchers to bridge gaps in SDDI.
Finally, two of our original themes, AI & SDDI and AI & CS Education, were cross-cutting themes for the four main themes, as they have potential implications for all other themes. For these, we discuss how the rapid developments in AI can challenge or support progress in SDDI. We also underline how a carefully balanced use of AI is essential for building an inclusive future that avoids undesirable scenarios and gets as close as possible to a desirable future.
@@ 2. Research Process
The results presented in this paper originate from facilitated discussions during an academic meeting in June 2023444https://shonan.nii.ac.jp/seminars/194. This meeting brought together 23 software engineering researchers and practitioners from diverse backgrounds and career phases interested in fostering SDDI-related research. 18 workshop participants were actively conducting SE research at academic institutions from around the world, while five were researchers and/or developers at companies in industry. The workshop was organized in a hybrid format, with seven participants joining remotely. The organizers of the meeting followed NII Shonan’s instructions for balanced participation, which are stated as follows: “Please make a good balance of international diversity, research fields, and a proportion of theoretical and practical areas. The list also should be a mixture of proven experts and promising young researchers, preferably including representative researchers of their community.”. All the participants, from PhD candidates to full professors, were experienced in working with SDDI topics.
Figure 1. Overview of our 1-2-4-ALL process for understanding opportunities for SDDI through SE Research
We utilized 1-2-4-ALL as a systematic data collection method to allow self-reflection and collaborative discussion, building toward consensus or shared understanding among participants and to foster inclusive participation and generate a wide range of ideas from the participants in the meeting (Lipmanowicz and McCandless, 2013) [^5] https://www.liberatingstructures.com/1-1-2-4-all/. An overview of our method is presented in Figure 1. The 1-2-4-ALL method is one of the Liberating structures methods, which are techniques to support lively discussions and foster engagement in a group setting. Other popular liberating structures methods are, e.g., Impromptu Networking and Crowd Sourcing (McCandless and Lipmanowicz, [n. d.]). Liberating structures can be referred to as a “destructive methodology”, as it tries to create a discussion and communication between participants without classic ways for one-way communication such as PowerPoint presentations (Ferguson et al., 2015).
For this study, we decided to use 1-2-4-ALL as it is a more structured way to approach shared decisions than, for example, the Impromptu Networking method, which is more spontaneous structure for sharing of ideas and engage the participants (Ferguson et al., 2015). Prior work has leveraged the 1-2-4-ALL method to explore techniques to improve student engagement in health and science education (Ferguson et al., 2015), identify research priorities in psychiatric-mental health (Mahoney et al., 2016), and enhance behavior and performance for management teams (Kjellström et al., 2022). In 1-2-4-ALL, ideation starts with an individual reflection (1 minute), followed by a discussion in pairs, building on ideas from self-reflection (2 minutes), sharing ideas in groups of four focusing on similarities and differences (4 minutes) and then sharing the best group idea with all the participants. In SE, liberating structures have been used especially as a discussion and decision tool in Agile teams or to ensure that everyone gets their voice heard in remote meetings (Turner, 2021; Kang et al., 2023).
@@
As the results of the final group discussion, the groups were asked to present one goal, two outcomes, and four themes related to challenges and opportunities to improve SDDI. The results of the breakout group discussions were shared with all participants to obtain feedback and reach an agreement on a defined list of SDDI themes. The participants were then divided into groups according to the defined themes to further expand on the research challenges and solutions. After intensive discussions around the identified goals, outcomes, and themes in multiple group sessions, we developed a report [^6] https://shonan.nii.ac.jp/docs/No%20.194.pdf summarizing the discussion for each theme.
@@
In this paper, we expand our results from the report to provide a more coherent vision of future research directions for SDDI. We expand the discussion around four of the main themes: methodologies and metrics, intersectionality, knowledge transfer, and connections to socioeconomic understanding. These themes were chosen based on the timeliness, potential impact, and interest participants attributed to the themes. From each of the themes, we identified potential benefits and harms. In turn, the themes, related research literature and the discussion in the SE 2030 workshop held at FSE Conference 2024, where we presented our preliminary results, guided us to form research goals for 2030 (Figure 2). In the discussion section, we discuss the two remaining themes presented in the original meeting: ‘AI & SDDI’ and ‘AI & CS Education’, as these themes were noticed to be cross-cutting.
Figure 2. Overview of the process for creating the SE 2030 Roadmap for SDDI. The process started from NII Shonan meeting, continued for the first version presented in SE 2030 Workshop and insights from there led to the final roadmap.
@@ 3. Four Scenarios for 2030
In the following, we present two desirable “utopian” scenarios, that is, scenarios “having the characteristics and organization of a perfect society” [^7] https://dictionary.cambridge.org/dictionary/english/utopian and two contrasting undesirable “dystopian” scenarios, i.e., scenarios “relating to a very bad or unfair society in which there is a lot of suffering” [^8] https://dictionary.cambridge.org/dictionary/english/dystopian. These scenarios capture the future of software development work and the future of education, two areas that we, as SE researchers, can actively study and shape. This approach of describing the potential impact of AI in the future in terms of desirable (“utopian”) and undesirable (“dystopian”) outcomes has been used in related work, e.g., (Aquino et al., 2023; Chan, 2023)
S1: Desirable Scenario in Workplace: Jamie works as a recently-graduated junior software engineer at an up-and-coming tech company. The company is at the forefront of inclusion, offering, e.g., flexible work arrangements that consider factors such as child care responsibilities, support for various career paths and the age of employees and applicants, and preferences concerning work and communication modes. These initiatives are paying off, as the company has successfully attracted a diverse workforce, leading to higher productivity, a welcoming and safe work environment, and more innovative and successful products. Despite coming from a minority group and being relatively junior, Jamie feels that the company provides a safe space for developing their career and competencies and invites them to participate in decision-making. For instance, the company offers automated tools to support various workflows, e.g., in code review and in programming, that allow Jamie to receive early feedback on their work, spot mistakes, and improve their skills without fear of discrimination. Similarly, the tools and practices at the company allow diverse teams to communicate effectively and positively. The company fosters a culture that encourages the active use and development of these tools for everyone’s benefit, both within and beyond the company. In addition, the company carefully monitors existing diversity metrics and plans interventions based on them. Regular surveys help company leadership to consider employees’ values in their strategic decisions. Finally, the company employs mechanisms at various stages to avoid potential misuse of data obtained from its employees, e.g., through transparency in decision-making.
@@
S2: Desirable Scenario in Education: Kris is a university-level educator in software engineering. Due to inclusive and equitable conditions at the university, students from varied backgrounds are enrolled in Kris’ courses, contributing to a truly diverse and inclusive study experience. Although classes are large so that the university can cope with the continuing strong demand for software engineering professionals, technology advances and government incentives make it possible for all students to receive targeted, individual feedback, regardless of diverse aspects such as gender, nationality, or age. In part, this is achieved through tools and indicators that enable automated feedback and coaching, so that Kris and their staff can fully focus on providing the best possible learning experience. Finally, the assessment has changed from standardized and unified exams and assignments to a personalized form, tailored to suit each student’s individual needs. In this environment, students feel safe to make mistakes and express themselves. In turn, this allows fruitful collaborations between students and staff.
To contrast the experiences of Jamie and Kris, we outline an undesirable future in which Ash struggles in their workplace, and Moss struggles in their role as a student.
S3: Undesirable Scenario in Workplace: Ash has graduated from a prestigious university in their home country with a degree in software engineering. They are employed at a local IT company. The company provides software development services to large corporations in rich and highly developed countries. Due to the economic realities in their home country, Ash and their company cannot afford to pay for existing top-of-the-shelf tools, such as coding coaches or LLM-based code generation tools. Even for free tools, unreliable and slow internet connections limit the use of these tools. This results in more manual work and less time to develop skills. Similarly, language barriers prevent them from learning and improving their skills at the same pace as similarly qualified graduates in richer and more developed countries, as existing tools cater only to English speakers. At Ash’s employer, traditional gender role models and hierarchical structures persist and affect career progression chances. As a result, Ash feels that their chance of a successful career is relatively low due to factors outside of their control. An overall lack of awareness of diversity matters also impedes the services provided by Ash’s employer, as stereotypes and biases are ingrained in the products provided by the company.
@@
S4: Undesirable Scenario in Education: Moss is a student in software engineering at a local university. Many tasks and activities that used to involve human interaction have been replaced by automated tools. This has made education cheaper for the university. However, instead of lowering tuition fees, student numbers have increased dramatically, but staff have been reduced. Lecture topics and assignments lack personalization and typically cover generic and stereotypical examples. As a result, Moss does not feel like they belong to the university, having almost no interaction with students and staff who come from a similar background as them. As someone who moved to the city for their studies, social integration has therefore been lacking. Instead, Moss feels that education could just as well be remote.
4. Results: Avoiding Undesirable Scenarios